# Large Language Models for Cyber Security: A Systematic Literature Review

### v3

@@ 3.1. Application of LLMs in Network Security
Web fuzzing. Web fuzzing is a mutation-based fuzzer that generates test cases incrementally based on the coverage feedback it receives from the instrumented web application (van Rooij et al., 2021). Security is undeniably the most critical concern for web applications. Fuzzing can help operators discover more potential security risks in web applications. Liang et al. (Liang et al., 2024) proposed GPTFuzzer based on an encoder-decoder architecture. It generates effective payloads for web application firewalls (WAFs) targeting SQL injection, XSS, and RCE attacks by generating fuzz test cases. The model undergoes reinforcement learning (Li et al., 2024) fine-tuning and KL-divergence penalty to effectively generate attack payloads and mitigate the local optimum issue. Similarly, Liu et al. (Liu et al., 2020a) utilized an encoder-decoder architecture model to generate SQL injection detection test cases for web applications, enabling the translation of user inputs into new test cases. Meng et al.’s CHATAFL (Meng et al., 2024), on the other hand, shifts focus to leveraging LLMs for generating structured and sequenced effective test inputs for network protocols lacking machine-readable versions.
@@
Penetration test. Conducting a controlled attack on a computer system to evaluate its security is the essence of penetration testing, which remains a pivotal approach utilized by organizations to bolster their defenses against cyber threats (Schwartz and Kurniawati, 2019). The general penetration testing process consists of three steps: information gathering, payload construction, and vulnerability exploitation. Temara (Temara, 2023) utilized LLMs to gather information for penetration testing, including the IP address, domain information, vendor technologies, SSL/TLS credentials, and other details of the target website. Sai Charan et al. (Charan et al., 2023) critically examined the capability of LLMs to generate malicious payloads for penetration testing, with results indicating that ChatGPT can generate more targeted and complex payloads for attackers. Happe et al. (Happe et al., 2023) developed an automated Linux privilege escalation guidance tool using LLMs. Additionally, the automated penetration testing tool PentestGPT (Deng et al., 2023a), based on LLMs, achieved excellent performance on a penetration testing benchmark containing 13 scenarios and 182 subtasks by combining three self-interacting modules (inference, generation, and parsing modules).
@@ 3.2. Application of LLMs in Software and System Security
Vulnerability detection. The escalation in software vulnerabilities is evident in the recent surge of vulnerability reports documented by Common Vulnerabilities and Exposures (CVEs) (Anon, 2022). With this rise, the potential for cybersecurity attacks grows, posing significant economic and social risks. Hence, the detection of vulnerabilities becomes imperative to safeguard software systems and uphold social and economic stability. The method of utilizing LLMs for static vulnerability detection in code shows significant performance improvements compared to traditional approaches based on graph neural networks or matching rules (Chen et al., 2023b; Ferrag et al., 2023; Bakhshandeh et al., 2023; Chen et al., 2023c; Wang et al., 2023c; Chow et al., 2023; Liu et al., 2023d; Thapa et al., 2022; Zhang et al., 2023f, g; Khare et al., 2023; Ullah et al., 2023; Quan et al., 2023). The potential demonstrated by GPT series models in vulnerability detection tasks is particularly evident (Liu et al., 2023d; Zhang et al., 2023f; Khare et al., 2023; Chen et al., 2023b; Ferrag et al., 2023; Ullah et al., 2024). However, LLMs may generate false positives when dealing with vulnerability detection tasks due to minor changes in function and variable names or modifications to library functions (Ullah et al., 2023). Liu et al. (Liu et al., 2023c) proposed LATTE, which combines LLMs to achieve automated binary taint analysis. This overcomes the limitations of traditional taint analysis, which requires manual customization of taint propagation rules and vulnerability inspection rules. They discovered 37 new vulnerabilities in real firmware. Tihanyi et al. (Tihanyi et al., 2023) used LLMs to generate a large-scale vulnerability-labeled dataset, FormAI, while also noting that over 50% of the code generated by LLMs may contain vulnerabilities, posing a significant risk to software security.
@@
Vulnerability repair. Due to the sharp increase in the number of detected vulnerabilities and the complexity of modern software systems, manually fixing security vulnerabilities is extremely time-consuming and labor-intensive for security experts (Zhang et al., 2022). Research shows that 50% of vulnerabilities have a lifecycle exceeding 438 days (Li and Paxson, 2017). Delayed vulnerability patching may result in ongoing attacks on software systems (Liu et al., 2020b), causing economic losses to users. The T5 model based on the encoder-decoder architecture performs better in vulnerability repair tasks (Fu et al., 2022; Zhang et al., 2023b). Although LLMs can effectively generate fixes, challenges remain in maintaining the functionality correctness of functions (Pearce et al., 2023), and they are susceptible to influences from different programming languages. For example, the current capabilities of LLMs in repairing Java vulnerabilities are limited (Wu et al., 2023a). Constructing a comprehensive vulnerability repair dataset and fine-tuning LLMs on it can significantly improve the model’s performance in vulnerability repair tasks (Fu et al., 2022). Alrashedy et al. (Charalambous et al., 2023) proposed an automated vulnerability repair tool driven by feedback from static analysis tools. Tol et al. (Tol and Sunar, 2023) proposed a method called ZeroLeak, which utilizes LLMs to repair side-channel vulnerabilities in programs. Charalambous et al. (Alrashedy and Aljasser, 2024) combined LLMs with Bounded Model Checking (BMC) to verify the effectiveness of repair solutions, addressing the problem of decreased functionality correctness after using LLMs to repair vulnerabilities.
@@
Bug detection. Bugs typically refer to any small faults or errors present in software or hardware, which may cause programs to malfunction or produce unexpected results. Some bugs may be exploited by attackers to create security vulnerabilities. Therefore, bug detection is crucial for the security of software and system. LLMs can be utilized to generate code lines and compare them with the original code to flag potential bugs within code snippets (Ahmad et al., 2023a). They can also combine feedback from static analysis tools to achieve precise bug localization (Jin et al., 2023; Li et al., 2023b). Fine-tuning techniques are crucial for bug detection tasks as well, applying fine-tuning allows LLMs to identify errors in code without relying on test cases (Lee et al., 2023; Yang et al., 2023c). Additionally, Du et al. (Du and Yu, 2023) and Li et al. (Li et al., 2023c) introduced the concept of contrastive learning, which focuses LLMs on the subtle differences between correct and buggy versions of code lines. Fang et al. (Fang et al., 2023) proposed a software-agnostic representation method called RepresentThemAll, based on contrastive learning and fine-tuning modules, suitable for various downstream tasks including bug detection and predicting the priority and severity of bugs.
@@
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023c), CodeT5 (Wang et al., 2023a; Tang et al., 2023; Huang et al., 2023), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023a), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023a), UniXcoder (Zhang et al., 2023c), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), and GPT Series (Zhang et al., 2023d; Olausson et al., 2024; Zhang et al., 2023c; Xia and Zhang, 2023b, a; Tang et al., 2023; Zhang et al., 2023h) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2023b). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023a). As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023a). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes. By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023b; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@
Program fuzzing. Fuzz testing, or fuzzing, refers to an automated testing method aimed at generating inputs to uncover unforeseen behaviors. Both researchers and practitioners have effectively developed practical fuzzing tools, demonstrating significant success in detecting numerous bugs and vulnerabilities within real-world systems (Boehme et al., 2021). The generation capability of LLMs enables testing against various input program languages and different features (Xia et al., 2024; Deng et al., 2023b), effectively overcoming the limitations of traditional fuzz testing methods. Under strategies such as repetitive querying, example querying, and iterative querying (Zhang et al., 2023a), LLMs can significantly enhance the generation effectiveness of test cases. LLMs can generate test cases that trigger vulnerabilities from historical bug reports of programs (Deng et al., 2023c), produce test cases similar but different from sample inputs (Hu et al., 2023c), analyze compiler source code to generate programs that trigger specific optimizations (Yang et al., 2023a), and split the testing requirements and test case generation using a dual-model interaction framework, assigning them to different LLMs for processing.
@@
Reverse engineering and binary analysis. Reverse engineering is the process of attempting to understand how existing artifacts work, whether for malicious purposes or defensive purposes, and it holds significant security implications. The capability of LLMs to recognize software functionality and extract important information enables them to perform certain reverse engineering steps (Pearce et al., 2022). For example, Xu et al. (Xu et al., 2023) achieved recovery of variable names from binary files by propagating LLMs query results through multiple rounds. Armengol-Estapé et al. (Armengol-Estapé et al., 2024) combined type inference engines with LLMs to perform disassembly of executable files and generate program source code. LLMs can also be used to assist in binary program analysis. Sun et al. (Sun et al., 2023a) proposed DexBert for characterizing Android system binary bytecode. Pei et al. (Pei et al., 2024) preserved the semantic symmetry of code based on group theory, resulting in their binary analysis framework SYMC demonstrating outstanding generalization and robustness in various binary analysis tasks. Song et al. (Song et al., 2022) utilized LLMs to address authorship analysis issues in software engineering, effectively applying them to real-world APT malicious software for organization-level verification. Some studies (Hu et al., 2024) apply LLMs to enhance the readability and usability of decompiler outputs, thereby assisting reverse engineers in better understanding binary files.
@@
Malware detection. Due to the rising volume and intricacy of malware, detecting malicious software has emerged as a significant concern. While conventional detection techniques rely on signatures and heuristics, they exhibit limited effectiveness against unknown attacks and are susceptible to evasion through obfuscation techniques (Bilot et al., 2023). LLMs can extract semantic features of malware, leading to more competitive performance. AVScan2Vec, proposed by Joyce et al. (Joyce et al., 2023), transforms antivirus scan reports into vector representations, effectively handling large-scale malware datasets and performing well in tasks such as malware classification, clustering, and nearest neighbor search. Botacin (Botacin, 2023) explored the application of LLMs in malware defense from the perspective of malware generation. While LLMs cannot directly generate complete malware based on simple instructions, they can generate building blocks of malware and successfully construct various malware variants by blending different functionalities and categories. This provides a new perspective for malware detection and defense.
@@
System log analysis. Analyzing the growing amount of log data generated by software-intensive systems manually is unfeasible due to its sheer volume. Numerous deep learning approaches have been suggested for detecting anomalies in log data. These approaches encounter various challenges, including dealing with high-dimensional and noisy log data, addressing class imbalances, and achieving generalization (Jakub and Branišová, 2017). Nowadays, researchers are utilizing the language understanding capabilities of LLMs to identify and analyze anomalies in log data. Compared to traditional deep learning methods, LLMs demonstrate outstanding performance and good interpretability (Qi et al., 2023a; Shan et al., 2024). Fine-tuning LLMs for specific types of logs (Karlsen et al., 2023) or using reinforcement learning-based fine-tuning strategies (Han et al., 2023) can significantly enhance their performance in log analysis tasks. LLMs are also being employed for log analysis in cloud servers (Liu et al., 2023b; Chen et al., 2023e), where their reasoning abilities can be combined with server logs to infer the root causes of cloud service incidents.

### v4

@@ 3.1. Application of LLMs in Network Security
Web fuzzing. Web fuzzing is a mutation-based fuzzer that generates test cases incrementally based on the coverage feedback it receives from the instrumented web application (van Rooij et al., 2021). Security is undeniably the most critical concern for web applications. Fuzzing can help operators discover more potential security risks in web applications. Liang et al. (Liang et al., 2024) proposed GPTFuzzer based on a decoder-only architecture. It generates effective payloads for web application firewalls (WAFs) targeting SQL injection, XSS, and RCE attacks by generating fuzz test cases. The model undergoes reinforcement learning (Li et al., 2024a) fine-tuning and KL-divergence penalty to effectively generate attack payloads and mitigate the local optimum issue. Similarly, Liu et al. (Liu et al., 2020a) utilized an encoder-decoder architecture model to generate SQL injection detection test cases for web applications, enabling the translation of user inputs into new test cases. Meng et al.’s CHATAFL (Meng et al., 2024a), on the other hand, shifts focus to leveraging LLMs for generating structured and sequenced effective test inputs for network protocols lacking machine-readable versions.
@@
Penetration test. Conducting a controlled attack on a computer system to evaluate its security is the essence of penetration testing, which remains a pivotal approach utilized by organizations to bolster their defenses against cyber threats (Schwartz and Kurniawati, 2019). The general penetration testing process consists of three steps: information gathering, payload construction, and vulnerability exploitation. Temara (Temara, 2023) utilized LLMs to gather information for penetration testing, including the IP address, domain information, vendor technologies, SSL/TLS credentials, and other details of the target website. Sai Charan et al. (Charan et al., 2023) critically examined the capability of LLMs to generate malicious payloads for penetration testing, with results indicating that ChatGPT can generate more targeted and complex payloads for attackers. Happe et al. (Happe et al., 2023) developed an automated Linux privilege escalation guidance tool using LLMs. Additionally, the automated penetration testing tool PentestGPT (Deng et al., 2024b), based on LLMs, achieved excellent performance on a penetration testing benchmark containing 13 scenarios and 182 subtasks by combining three self-interacting modules (inference, generation, and parsing modules).
Protocol analysis. LLMs are being applied to analyze complex network protocol specifications for detecting inconsistencies. For example, CellularLint utilizes domain-adapted LLMs combined with few-shot and active learning techniques to systematically identify behavioral inconsistencies in 4G/5G cellular standards (Rahman et al., 2024). Such inconsistencies found in specifications can impact the security and interoperability of network implementations.
Protocol fuzzing. LLMs are utilized to guide protocol fuzzing, particularly for protocols specified only in natural language, overcoming limitations of traditional fuzzers in generating structurally and sequentially valid inputs. Approaches like ChatAFL integrate LLMs into the fuzzing loop, using the model’s understanding of message types and sequences to generate valid test inputs and guide mutation-based fuzzers like AFL++ (Meng et al., 2024b). Other frameworks, such as LLMIF, augment LLMs with protocol specifications to automatically extract message formats, field constraints, and state transition rules, and to reason about device responses during fuzzing, specifically targeting IoT protocols (Wang et al., 2024d). These LLM-guided approaches have demonstrated significant improvements in code coverage and vulnerability discovery compared to baseline fuzzers.
Web vulnerability detection. LLMs are also employed for detecting vulnerabilities in web applications, although challenges exist regarding language-specific characteristics and data availability. Frameworks like RealVul aim to enhance the performance of LLMs for detecting PHP vulnerabilities (Cao et al., 2024). This is achieved by integrating deep program analysis techniques, such as control and data flow analysis, to identify vulnerability candidates and extract relevant code slices.
@@ 3.2. Application of LLMs in Software and System Security
Vulnerability detection. The escalation in software vulnerabilities is evident in the recent surge of vulnerability reports documented by Common Vulnerabilities and Exposures (CVEs) (Anon, 2022). With this rise, the potential for cybersecurity attacks grows, posing significant economic and social risks. Hence, the detection of vulnerabilities becomes imperative to safeguard software systems and uphold social and economic stability. The method of utilizing LLMs for static vulnerability detection in code shows significant performance improvements compared to traditional approaches based on graph neural networks or matching rules (Chen et al., 2023b; Ferrag et al., 2023; Bakhshandeh et al., 2023; Chen et al., 2023c; Wang et al., 2023d; Chow et al., 2023; Liu et al., 2023d; Thapa et al., 2022; Zhang et al., 2024b, 2023e; Khare et al., 2023; Ullah et al., 2023; Quan et al., 2023). For instance, SCALE enhances detection by using LLMs to generate comments for Abstract Syntax Tree nodes, creating a novel Structured Natural Language Comment Tree (SCT) representation fed to a graph classifier (Wen et al., 2024). Furthermore, techniques like multi-task instruction fine-tuning, as used in VulLLM, aim to improve model generalization by training simultaneously on detection, localization, and vulnerability explanation tasks (Du et al., 2024a). The potential demonstrated by GPT series models in vulnerability detection tasks is particularly evident (Liu et al., 2023d; Zhang et al., 2024b; Khare et al., 2023; Chen et al., 2023b; Ferrag et al., 2023; Ullah et al., 2024; Lin and Mohaisen, 2025; Wu et al., 2024b; Xu et al., 2024a). However, LLMs may generate false positives when dealing with vulnerability detection tasks due to minor changes in function and variable names or modifications to library functions (Ullah et al., 2023). Liu et al. (Liu et al., 2023c) proposed LATTE, which combines LLMs to achieve automated binary taint analysis. This overcomes the limitations of traditional taint analysis, which requires manual customization of taint propagation rules and vulnerability inspection rules. They discovered 37 new vulnerabilities in real firmware. Tihanyi et al. (Tihanyi et al., 2023) used LLMs to generate a large-scale vulnerability-labeled dataset, FormAI, while also noting that over 50% of the code generated by LLMs may contain vulnerabilities, posing a significant risk to software security.
@@
Vulnerability repair. Due to the sharp increase in the number of detected vulnerabilities and the complexity of modern software systems, manually fixing security vulnerabilities is extremely time-consuming and labor-intensive for security experts (Zhang et al., 2022). Research shows that 50% of vulnerabilities have a lifecycle exceeding 438 days (Li and Paxson, 2017). Delayed vulnerability patching may result in ongoing attacks on software systems (Liu et al., 2020b), causing economic losses to users. The T5 model based on the encoder-decoder architecture performs better in vulnerability repair tasks (Fu et al., 2022; Zhang et al., 2023a). Although LLMs can effectively generate fixes, challenges remain in maintaining the functionality correctness of functions (Pearce et al., 2023), and they are susceptible to influences from different programming languages (Le et al., 2024). For example, the current capabilities of LLMs in repairing Java vulnerabilities are limited (Wu et al., 2023a), while studies also investigate their effectiveness on JavaScript vulnerabilities, showing performance improvements with better contextual prompts (Le et al., 2024). Constructing a comprehensive vulnerability repair dataset and fine-tuning LLMs on it can significantly improve the model’s performance in vulnerability repair tasks (Fu et al., 2022; Li et al., 2024b). To assist developers, especially when vulnerable code locations are untracked in issue reports, approaches like PATUNTRACK use LLMs with Retrieval-Augmented Generation (RAG) and few-shot learning to generate relevant patch examples (insecure code, patch, explanation) (Jiang et al., 2024). Another direction involves generating natural language repair suggestions instead of direct code patches; VulAdvisor first uses LLMs to create a dataset of suggestions from code/patch pairs and then fine-tunes CodeT5 with context-aware attention to generate such suggestions directly from vulnerable code (Zhang et al., 2024d). Alrashedy et al. (Charalambous et al., 2023) proposed an automated vulnerability repair tool driven by feedback from static analysis tools. Tol et al. (Tol and Sunar, 2024) proposed a method called ZeroLeak, which utilizes LLMs to repair side-channel vulnerabilities in programs. Charalambous et al. (Alrashedy and Aljasser, 2024) combined LLMs with Bounded Model Checking (BMC) to verify the effectiveness of repair solutions, addressing the problem of decreased functionality correctness after using LLMs to repair vulnerabilities (Chen and Jabbarvand, 2024).
@@
Bug detection. Bugs typically refer to any small faults or errors present in software or hardware, which may cause programs to malfunction or produce unexpected results. Some bugs may be exploited by attackers to create security vulnerabilities. Therefore, bug detection is crucial for the security of software and system. LLMs can be utilized to generate code lines and compare them with the original code to flag potential bugs within code snippets (Ahmad et al., 2023). They can also combine feedback from static analysis tools to achieve precise bug localization (Jin et al., 2023; Li et al., 2023b; Hossain et al., 2024). Fine-tuning techniques are crucial for bug detection tasks as well, applying fine-tuning allows LLMs to identify errors in code without relying on test cases (Lee et al., 2023; Yang et al., 2023b; Cai et al., 2024b). To address the challenge of hallucinations or false positives in LLM-generated bug reports, frameworks like LLMSAN use techniques such as Chain-of-Thought prompting combined with post-hoc validation of the LLM’s claimed evidence (e.g., data-flow paths) via program analysis and targeted queries (Wang et al., 2024e). Additionally, Du et al. (Du and Yu, 2023) and Li et al. (Li et al., 2023c) introduced the concept of contrastive learning, which focuses LLMs on the subtle differences between correct and buggy versions of code lines. Fang et al. (Fang et al., 2023) proposed a software-agnostic representation method called RepresentThemAll, based on contrastive learning and fine-tuning modules, suitable for various downstream tasks including bug detection and predicting the priority and severity of bugs.
@@
Bug repair. LLMs possess robust code generation capabilities, and their utilization in engineering for code generation can significantly enhance efficiency. However, code produced by LLMs often carries increased security risks, such as bugs and vulnerabilities (Perry et al., 2023). These program bugs can lead to persistent security vulnerabilities. Hence, automating the process of bug fixing is imperative, involving the use of automation technology to analyze flawed code and generate accurate patches to rectify identified issues. LLMs like CodeBERT (Le-Cong et al., 2023; Xia and Zhang, 2022; Huang et al., 2023; Zhang et al., 2023b), CodeT5 (Wang et al., 2023b; Tang et al., 2023; Huang et al., 2023; Liu et al., 2024c), Codex (Jin et al., 2023; Fan et al., 2023; Xia and Zhang, 2023; Chow et al., 2024; Tang et al., 2024), LLaMa (Tang et al., 2023), CodeLLaMa (Olausson et al., 2024; Silva et al., 2024) , CodeGEN (Xia and Zhang, 2023), UniXcoder (Zhang et al., 2023b), T5 (Yuan et al., 2022), PLBART (Huang et al., 2023), Starcoder (Zhao et al., 2024) and GPT Series (Zhang et al., 2023c; Olausson et al., 2024; Zhang et al., 2023b; Xia and Zhang, 2024, 2023; Tang et al., 2023; Zhang et al., 2023f; Yang et al., 2024b; Yin et al., 2024b; Zhang et al., 2024c; Huang et al., 2024; Feng and Chen, 2024; Zhong et al., 2024; Wang et al., 2024c) have showcased effectiveness in generating syntactically accurate and contextually relevant code. This includes frameworks with encoder-decoder architecture like Repilot (Wei et al., 2023c), tailored specifically for producing repair patches. Utilizing LLMs for program repair can achieve competitive performance in producing patches for various types of errors and defects (Xia and Zhang, 2024). These models effectively capture the underlying semantics and dependencies in code, resulting in precise and efficient patches. Moreover, fine-tuning LLMs on specific code repair datasets can further improve their ability to generate high-quality patches for real-world software projects. Integrating LLMs into program repair not only speeds up the error-fixing process but also allows software developers to focus on more complex tasks, thereby enhancing the reliability and maintainability of the software (Xia and Zhang, 2023). As demonstrated in the case of ChatGPT, notably enhances the accuracy of program repairs when integrated with interactive feedback loops (Xia and Zhang, 2023). This iterative process of patch generation and validation fosters a nuanced comprehension of software semantics, thereby resulting in more impactful fixes. Complementary to generating fixes, evaluating the capabilities of LLM-driven code agents requires realistic benchmarks; SWT-Bench, for example, provides a framework based on real-world data to jointly assess test generation and fix validation performance (Mündler et al., 2024). By integrating domain-specific knowledge and technologies with the capabilities of LLMs, their performance is further enhanced. Custom prompts, fine-tuning for specific tasks, retrieving external data, and utilizing static analysis tools (Xia et al., 2023; Tang et al., 2023; Zhang et al., 2023a; Fu et al., 2022; Jin et al., 2023) significantly improve the effectiveness of bug fixes driven by LLMs.
@@
Program fuzzing. Fuzz testing, or fuzzing, refers to an automated testing method aimed at generating inputs to uncover unforeseen behaviors. Both researchers and practitioners have effectively developed practical fuzzing tools, demonstrating significant success in detecting numerous bugs and vulnerabilities within real-world systems (Boehme et al., 2021). The generation capability of LLMs enables testing against various input program languages and different features (Xia et al., 2024; Deng et al., 2023; Fu et al., 2024), effectively overcoming the limitations of traditional fuzz testing methods. Under strategies such as repetitive querying, example querying, and iterative querying (Zhang et al., 2024e), LLMs can significantly enhance the generation effectiveness of test cases. LLMs can generate test cases that trigger vulnerabilities from historical bug reports of programs (Deng et al., 2024c), produce test cases similar but different from sample inputs (Hu et al., 2023b), analyze compiler source code to generate programs that trigger specific optimizations (Yang et al., 2024a), and split the testing requirements and test case generation using a dual-model interaction framework, assigning them to different LLMs for processing. LLMs are also used to analyze program documentation to predict high-risk option combinations for targeted fuzzing, as demonstrated by ProphetFuzz (Wang et al., 2024f). Furthermore, LLMs facilitate the automatic generation of fuzz drivers for libraries (Lyu et al., 2024; Zhang et al., 2024f; Xu et al., 2024b), for instance, PROMPTFUZZ uses an iterative prompt fuzzing loop guided by code coverage feedback. To improve mutation effectiveness, especially for complex inputs like JavaScript, techniques like CovRL integrate coverage-guided reinforcement learning to fine-tune LLM-based mutators (Eom et al., 2024).
@@
Reverse engineering and binary analysis. Reverse engineering is the process of attempting to understand how existing artifacts work, whether for malicious purposes or defensive purposes, and it holds significant security implications. The capability of LLMs to recognize software functionality and extract important information enables them to perform certain reverse engineering steps (Pearce et al., 2022). For example, Xu et al. (Xu et al., 2023) achieved recovery of variable names from binary files by propagating LLMs query results through multiple rounds. Techniques like GENNM further advance this by using generative LLMs with context-aware fine-tuning, preference optimization, and iterative inference to recover names from stripped binaries (Xu et al., 2025). Armengol-Estapé et al. (Armengol-Estapé et al., 2024) combined type inference engines with LLMs to perform disassembly of executable files and generate program source code. LLMs can also be used to assist in binary program analysis. Sun et al. (Sun et al., 2023) proposed DexBert for characterizing Android system binary bytecode. Pei et al. (Pei et al., 2024) preserved the semantic symmetry of code based on group theory, resulting in their binary analysis framework SYMC demonstrating outstanding generalization and robustness in various binary analysis tasks. Song et al. (Song et al., 2022) utilized LLMs to address authorship analysis issues in software engineering, effectively applying them to real-world APT malicious software for organization-level verification. Some studies (Hu et al., 2024a; She et al., 2024; Fang et al., 2024b) apply LLMs to enhance the readability and usability of decompiler outputs, including decompiling WebAssembly into higher-level languages like C/C++, thereby assisting reverse engineers in better understanding binary files.
@@
Malware detection. Due to the rising volume and intricacy of malware, detecting malicious software has emerged as a significant concern. While conventional detection techniques rely on signatures and heuristics, they exhibit limited effectiveness against unknown attacks and are susceptible to evasion through obfuscation techniques (Bilot et al., 2023). LLMs can extract semantic features of malware, leading to more competitive performance. AVScan2Vec, proposed by Joyce et al. (Joyce et al., 2023), transforms antivirus scan reports into vector representations, effectively handling large-scale malware datasets and performing well in tasks such as malware classification, clustering, and nearest neighbor search. To address dataset limitations in specific domains like malicious package detection, approaches like Maltracker combine fine-grained code analysis with LLM-generated synthetic malicious samples to create enhanced datasets for training detection models  (Yu et al., 2024). Botacin (Botacin, 2023a) explored the application of LLMs in malware defense from the perspective of malware generation. While LLMs cannot directly generate complete malware based on simple instructions, they can generate building blocks of malware and successfully construct various malware variants by blending different functionalities and categories. This provides a new perspective for malware detection and defense.
@@
System log analysis. Analyzing the growing amount of log data generated by software-intensive systems manually is unfeasible due to its sheer volume. Numerous deep learning approaches have been suggested for detecting anomalies in log data. These approaches encounter various challenges, including dealing with high-dimensional and noisy log data, addressing class imbalances, and achieving generalization (Jakub and Branišová, 2017). Nowadays, researchers are utilizing the language understanding capabilities of LLMs to identify and analyze anomalies in log data. Studies compare different strategies, such as supervised fine-tuning (SFT) of models versus few-shot in-context learning (ICL) with LLMs, finding SFT often yields higher accuracy given sufficient labeled data, while ICL is advantageous for rapid deployment or limited data scenarios (Jin et al., 2024). Compared to traditional deep learning methods, LLMs demonstrate outstanding performance and good interpretability (Qi et al., 2023a; Shan et al., 2024). Fine-tuning LLMs for specific types of logs (Karlsen et al., 2023; Jeong et al., 2025), using reinforcement learning-based fine-tuning strategies (Han et al., 2023) can significantly enhance their performance in log analysis tasks. LLMs are also being employed for log analysis in cloud servers (Liu et al., 2023b; Chen et al., 2023d), where their reasoning abilities can be combined with server logs to infer the root causes of cloud service incidents.
Secure code generation. While LLMs demonstrate strong code generation capabilities, the security of the generated code is a significant concern, as models trained on large corpora may replicate insecure patterns (Perry et al., 2023). Studies investigating LLM-generated code that utilizes security APIs, such as Java security APIs, have found high rates of misuse, highlighting the risks of relying on LLMs for security-sensitive implementations (Mousavi et al., 2024). To mitigate these risks, research explores methods to guide LLMs towards more secure code generation; for example, PromSec proposes an interactive loop involving static analysis, graph-based code fixing, and reverse-prompt engineering to automatically optimize prompts for generating functionally correct and more secure code (Nazzal et al., 2024).
ADS security verification. LLMs are also being applied to the safety and security verification of complex software-intensive systems like Autonomous Driving Systems (ADS). One application involves leveraging LLMs to enhance simulation testing; for instance, SoVAR utilizes LLMs to automatically extract and generalize test scenarios from natural language accident reports, generating more diverse and realistic test cases for simulation (Guo et al., 2024). Another application focuses on automating the analysis of test results; DIAVIO employs fine-tuned LLMs, informed by real-world accident data represented in a domain-specific language, to automatically diagnose the cause and type of safety violations identified during ADS simulation testing (Lu et al., 2024).
Code obfuscation. Analyzing obfuscated code is crucial for tasks like malware analysis and understanding protected software, and researchers have begun evaluating LLMs for this purpose. Systematic studies assessing models like the GPT and LLaMA families find that while larger models perform well on non-obfuscated code, their analysis accuracy significantly decreases when dealing with obfuscated code, and they exhibit limited capability in generating de-obfuscated code (Fang et al., 2024a).